3.10 List of Mathematical Symbols

In the following tables you find all the symbols normally accessible from math mode.

To use the symbols listed in Tables 3.12–3.16^{5}, the package amssymb must be loaded in the preamble of the document and the AMS math fonts must be installed on the system. If the AMS package and fonts are not installed, on your system, have a look at

CTAN:/tex-archive/macros/latex/packages/amslatex

Table 3.1: Math Mode Accents

 â \hat{a}     ǎ \check{a}   ã \tilde{a}     á \acute{a}
 à \grave{a}   ȧ \dot{a}     ä \ddot{a}      ă \breve{a}
 ā \bar{a}     a⃗ \vec{a}     Â \widehat{A}   Ã \widetilde{A}

Table 3.2: Lowercase Greek Letters

   α \alpha  θ \theta  o o  υ \upsilon
   β \beta  ϑ \vartheta π \pi  ϕ \phi
   γ \gamma  ι \iota  ϖ \varpi  φ \varphi
   δ \delta  κ \kappa  ρ \rho  χ \chi
   ϵ \epsilon  λ \lambda  ϱ \varrho  ψ \psi
   ε \varepsilon μ \mu  σ \sigma  ω \omega
   ζ \zeta  ν \nu  ς \varsigma
   η \eta  ξ \xi  τ \tau

Table 3.3: Uppercase Greek Letters

        Γ \Gamma Λ \Lambda Σ \Sigma  Ψ \Psi
        Δ \Delta Ξ \Xi  Υ \Upsilon Ω \Omega
        Θ \Theta Π \Pi  Φ \Phi

5These tables were derived from symbols.tex by David Carlisle and subsequently changed extensively as suggested by Josef Tkadlec

Table 3.4: Binary Relations

You can produce corresponding negations by adding a \not command as prefix to the following symbols.

< <  > >  = =
≤ \leq or \le  ≥ \geq or \ge  ≡ \equiv
≪ \ll  ≫ \gg           ≐ \doteq
≺ \prec  ≻ \succ  ∼ \sim
⪯ \preceq  ⪰ \succeq  ≃ \simeq
⊂ \subset  ⊃ \supset  ≈ \approx
⊆ \subseteq  ⊇ \supseteq     ≅ \cong
⊏ \sqsubset ^{a} ⊐ \sqsupset ^{a} ⋈ \Join ^{a}
⊑ \sqsubseteq ⊒ \sqsupseteq ▷◁ \bowtie
∈ \in  ∋ \ni , \owns  ∝ \propto
⊢ \vdash  ⊣ \dashv  ⊧ \models
| \mid  ∥ \parallel  ⊥ \perp
⌣ \smile  ⌢ \frown  ≍ \asymp
: :             ∉ \notin  ≠ \neq or \ne

aUse the latexsym package to access this symbol

Table 3.5: Binary Operators

+ +  − -
± \pm  ∓ \mp  ◁ \triangleleft
· \cdot  ÷ \div  ▷ \triangleright
× \times  \ \setminus  ⋆ \star
∪ \cup  ∩ \cap  ∗ \ast
⊔ \sqcup  ⊓ \sqcap  ∘ \circ
∨ \vee , \lor  ∧ \wedge , \land  • \bullet
⊕ \oplus  ⊖ \ominus  ⋄ \diamond
⊙ \odot  ⊘ \oslash  ⊎ \uplus
⊗ \otimes  ◯ \bigcirc  ⨿ \amalg
△ \bigtriangleup ▽ \bigtriangledown † \dagger
⊲ \lhd ^{a}  ⊳ \rhd ^{a}  ‡ \ddagger
⊴ \unlhd ^{a}  ⊵ \unrhd ^{a}  ≀ \wr

44  Typesetting Mathematical Formulae

Table 3.6: BIG Operators

     ∑ \sum  ⋃ \bigcup  ⋁ \bigvee  ⨁ \bigoplus
     ∏ \prod  ⋂ \bigcap  ⋀ \bigwedge ⨂ \bigotimes
     ∐ \coprod ⨆ \bigsqcup  ⨀ \bigodot
     ∫ \int  ∮ \oint  ⨄ \biguplus

Table 3.7: Arrows

← \leftarrow or \gets ⟵ \longleftarrow  ↑ \uparrow
→ \rightarrow or \to  ⟶ \longrightarrow  ↓ \downarrow
↔ \leftrightarrow  ⟷ \longleftrightarrow ↕ \updownarrow
⇐ \Leftarrow  ⟸ \Longleftarrow  ⇑ \Uparrow
⇒ \Rightarrow  ⟹ \Longrightarrow  ⇓ \Downarrow
⇔ \Leftrightarrow  ⟺ \Longleftrightarrow ⇕ \Updownarrow
↦ \mapsto  ⟼ \longmapsto  ↗ \nearrow
↩ \hookleftarrow  ↪ \hookrightarrow  ↘ \searrow
↼ \leftharpoonup  ⇀ \rightharpoonup  ↙ \swarrow
↽ \leftharpoondown  ⇁ \rightharpoondown  ↖ \nwarrow
⇌ \rightleftharpoons ⟺ \iff (bigger spaces)  ⇝ \leadsto ^{a}

aUse the latexsym package to access this symbol

Table 3.8: Delimiters

 ( (  ) )  ↑ \uparrow  ⇑ \Uparrow
 [ [ or \lbrack ] ] or \rbrack ↓ \downarrow  ⇓ \Downarrow
 { \{ or \lbrace } \} or \rbrace ↕ \updownarrow ⇕ \Updownarrow
 ⟨ \langle  ⟩ \rangle  | | or \vert  ∥ \| or \Vert
 ⌊ \lfloor  ⌋ \rfloor  ⌈ \lceil  ⌉ \rceil

/ /  \ \backslash  . (dual. empty)

Table 3.9: Large Delimiters

   { \lgroup     } \rgroup     { \lmoustache  } \rmoustache
    | \arrowvert ∥ \Arrowvert  | \bracevert

3.10 List of Mathematical Symbols  45

Table 3.10: Miscellaneous Symbols

. . . \dots  · · · \cdots       ⋮ \vdots  ... \ddots
ℏ \hbar  ı \imath  ȷ \jmath  ℓ \ell
ℜ \Re  ℑ \Im  ℵ \aleph  ℘ \wp
∀ \forall  ∃ \exists  ℧ \mho ^{a}  ∂ \partial
′ '  ′ \prime  ∅ \emptyset ∞ \infty
∇ \nabla  △ \triangle  □ \Box ^{a}  ◇ \Diamond ^{a}
⊥ \bot  ⊤ \top  ∠ \angle  √ \surd
♢ \diamondsuit ♡ \heartsuit ♣ \clubsuit ♠ \spadesuit
¬ \neg or \lnot ♭ \flat  ♮ \natural  ♯ \sharp

aUse the latexsym package to access this symbol

Table 3.11: Non-Mathematical Symbols

These symbols can also be used in text mode.

                † \dag  § \S   \begin{matrix}c \\ ◯ \end{matrix} \copyright
                ‡ \ddag ¶ \P £ \pounds

Table 3.12: AMS Delimiters

    ⌜ \ulcorner ⌝ \urcorner ⌞ \llcorner ⌟ \lrcorner

Table 3.13: AMS Greek and Hebrew

 𝟋 \digamma κ \varkappa ℶ \beth ℸ \daleth ℷ \gimel

Table 3.14: AMS Binary Relations

⋖ \lessdot  ⋗ \gtrdot  ≑ \doteqdot or \Doteq
⩽ \leqslant  ⩾ \geqslant  ≓ \risingdotseq
⪕ \eqslantless  ⪖ \eqslantgtr  ≒ \fallingdotseq
≦ \leqq  ≧ \geqq  ≖ \eqcirc
≪ \lll or \llless  ≫ \ggg or \gggtr  ≗ \circeq
≲ \lesssim  ≳ \gtrsim  ≜ \triangleq
⪅ \lessapprox  ⪆ \gtrapprox  ≏ \bumpeq
≶ \lessgtr  ≷ \gtrless  ≎ \Bumpeq
⋚ \lesseqgtr  ⋛ \gtreqless  ∼ \thicksim
⪋ \lesseqqgtr  ⪌ \gtreqqless  ≈ \thickapprox
≼ \preccurlyeq  ≽ \succcurlyeq  ≊ \approxeq
⋞ \curlyeqprec  ⋟ \curlyeqsucc  ∽ \backsim
≾ \precsim  ≿ \succsim  ⋍ \backsimeq
≾ \precapprox  ≿ \succapprox  ⊨ \vDash
⫅ \subseteqq  ⫆ \supseteqq  ⊩ \Vdash
⋐ \Subset  ⋑ \Supset  ⊪ \Vvdash
⊏ \sqsubset  ⊐ \sqsupset  ϶ \backepsilon
∴ \therefore  ∵ \because  ∝ \varpropto
 ∣ \shortmid  ∥ \shortparallel  ≬ \between
⌣ \smallsmile  ⌢ \smallfrown  ⋔ \pitchfork
◁ \vartriangleleft ▷ \vartriangleright ◀ \blacktriangleleft
⊴ \trianglelefteq  ⊵ \trianglerighteq  ▶ \blacktriangleright

Table 3.15: AMS Arrows

←−− \dashleftarrow  −−→ \dashrightarrow  ⊸ \multimap
⇇ \leftleftarrows  ⇉ \rightrightarrows  ⇈ \upuparrows
⇆ \leftrightarrows  ⇄ \rightleftarrows  ⇊ \downdownarrows
⇚ \Lleftarrow  ⇛ \Rrightarrow  ↿ \upharpoonleft
↞ \twoheadleftarrow  ↠ \twoheadrightarrow  ↾ \upharpoonright
↢ \leftarrowtail  ↣ \rightarrowtail  ⇃ \downharpoonleft
⇋ \leftrightharpoons ⇌ \rightleftharpoons  ⇂ \downharpoonright
↰ \Lsh  ↱ \Rsh  ⇝ \rightsquigarrow
↫ \looparrowleft  ↬ \looparrowright  ↭ \leftrightsquigarrow
↶ \curvearrowleft  ↷ \curvearrowright
⟲ \circlearrowleft  ⟳ \circlearrowright

Table 3.16: AMS Negated Binary Relations and Arrows

≮ \nless  ≯ \ngtr  ⫋ \varsubsetneqq
⪇ \lneq  ⪈ \gneq  ⫌ \varsupsetneqq
≰ \nleq  ≱ \ngeq  ⫅̸ \nsubseteqq
⩽̸ \nleqslant  ⩾̸ \ngeqslant  ⫆̸ \nsupseteqq
≨ \lneqq  ≩ \gneqq  ∤ \nmid
≨ \lvertneqq  ≩ \gvertneqq  ∦ \nparallel
≦̸ \nleqq  ≧̸ \ngeqq  ∤ \nshortmid
⋦ \lnsim  ⋧ \gnsim  ∦ \nshortparallel
⪉ \lnapprox  ⪊ \gnapprox  ≁ \nsim
⊀ \nprec  ⊁ \nsucc  ≇ \ncong
⪯̸ \npreceq  ⪰̸ \nsucceq  ⊬ \nvdash
⪵ \precneqq  ⪶ \succneqq  ⊭ \nvDash
⋨ \precnsim  ⋩ \succnsim  ⊮ \nVdash
⪹ \precnapprox  ⪺ \succnapprox  ⊯ \nVDash
⊊ \subsetneq  ⊋ \supsetneq  ⋪ \ntriangleleft
⊊ \varsubsetneq ⊋ \varsupsetneq ⋫ \ntriangleright
⊈ \nsubseteq  ⊉ \nsupseteq  ⋬ \ntrianglelefteq
⫋ \subsetneqq  ⫌ \supsetneqq  ⋭ \ntrianglerighteq
↚ \nleftarrow  ↛ \nrightarrow  ↮ \nleftrightarrow
⇍ \nLeftarrow  ⇏ \nRightarrow  ⇎ \nLeftrightarrow

Table 3.17: AMS Binary Operators

∔ \dotplus  ▪ \centerdot  ⊺ \intercal
⋉ \ltimes  ⋊ \rtimes  ⋇ \divideontimes
⋓ \Cup or \doublecup ⋒ \Cap or \doublecap ∖ \smallsetminus
⊻ \veebar  ⊼ \barwedge  ⩞ \doublebarwedge
⊞ \boxplus  ⊟ \boxminus  ⊝ \circleddash
⊠ \boxtimes  ⊡ \boxdot  ⊚ \circledcirc
⋋ \leftthreetimes  ⋌ \rightthreetimes ⊛ \circledast
⋎ \curlyvee  ⋏ \curlywedge

Table 3.18: AMS Miscellaneous

ℏ \hbar  ℏ \hslash  𝕜 \Bbbk
□ \square  ■ \blacksquare  Ⓢ \circledS
△ \vartriangle  ▲ \blacktriangle  ∁ \complement
▽ \triangledown ▼ \blacktriangledown ⅁ \Game
♢ \lozenge  ♦ \blacklozenge  ⋆ \bigstar
∠ \angle  ∡ \measuredangle  ∢ \sphericalangle
⧸ \diagup  ⧹ \diagdown  ‵ \backprime
∄ \nexists  Ⅎ \Finv  ∅ \varnothing
ð \eth  ℧ \mho

Table 3.19: Math Alphabets

Example Command  Required package

ABCdef \mathrm{ABCdef}
ABCdef \mathit{ABCdef}
ABCdef \mathnormal{ABCdef}
ABC  \mathcal{ABC}

𝒜ℬ𝒞  \mathcal{ABC}  eucal with option: mathcal
or  \mathscr{ABC}  eucal with option: mathscr

𝔄𝔅ℭ𝔡𝔢𝔣 \mathfrak{ABCdef}  eufrak

𝔸𝔹ℂ  \mathbb{ABC}  amsfonts or amssymb

Table 4.1: Fonts

\textrm{...} roman  \textsf{...} sans serif
\texttt{...}  typewriter
\textmd{...} medium  \textbf{...} bold face
\textup{...} upright  \textit{...} italic
\textsl{...} slanted  \textsc{...} small caps
  \emph{...} emphasised  \textnormal{...} document font

Table 4.2: Font sizes

\tiny  tiny font
\scriptsize  very small font
\footnotesize quite small font
\small  small font

\normalsize  normal font

\large  large font

\Large larger font
\LARGE very large font

                 \huge huge

\Huge largest

Table 4.3: Math fonts

Command  Example  Output

\mathcal{...}     $\mathcal{B}=c$  B = c
\mathrm{...}      $\mathrm{K}_2$  K_{2}
\mathbf{...}      $\sum x=\mathbf{v}$  ∑ x = v
\mathsf{...}      $\mathsf{G\times R}$  G × R
\mathtt{...}      $\mathtt{L}(b,c)$                 L(b, c)
\mathnormal{...}  $\mathnormal{R_{19}}\neq R_{19}$ R_{19} ≠ R_{19}
\mathit{...}      $\mathit{ffi}\neq ffi$  ffi ≠ ffi